OpenStudy (anonymous):

Use Property 1 for radicals to simplify the following radical expression as much as possible. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }\sqrt{24x^3}\]

hartnn (hartnn):

can you factor 24 ?

OpenStudy (anonymous):

4 & 6

hartnn (hartnn):

and \(x^3 = x \times x^2\) so, \(\Large \sqrt {24x^3 } = \sqrt 4 \sqrt 6 \sqrt x \sqrt {x^2} \)

hartnn (hartnn):

simplify 1st and last term

OpenStudy (anonymous):

2 & x

hartnn (hartnn):

\(\Large \sqrt {24x^3 } = \sqrt 4 \sqrt 6 \sqrt x \sqrt {x^2} = 2x \sqrt {6x}\) thats it! cannot be further simplified.

OpenStudy (anonymous):

Yaaay! Thanks!

hartnn (hartnn):

don't forget the initial 3/2

hartnn (hartnn):

so final answer will be \(3x \sqrt {6x}\)